OpenStudy (anonymous):

Ken purchases 5 boxes of flooring and 3 pints of sealant for $70.75. Jenna buys 6 boxes of flooring and 4 pints of sealant for $86.00. How much more does a box of flooring cost than a pint of sealant?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{5b+3p=70.75\\~\\ 6b+4p=86\\~\\ }\end{align}\) solve both equations find \(\large b\) and \(\large p\) \(\large b-p\) will be the answer

OpenStudy (anonymous):

How do I find b and P ?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ 6b+4p=86--A\\~\\ 3b+2p=43\\~\\ 3b=43-2p\\~\\ \bf b=\dfrac{43-2p}{3}\\~\\ \tt 5b+3p=70.75--B\\~\\ \tt 5(\dfrac{43-2p}{3})+3p=70.75 }\end{align}\)

OpenStudy (mathmath333):

now solve for \(\Large p\)

OpenStudy (mathmath333):

then after finding \(p\) put its value in \(\Large A\) and find \(b\)